Your firm designs training materials for computer training classes, and you have just received a request to bid on a contract to produce a complete set of training manuals for an 8-session class. From previous experience, you know that your firm follows an 85% learning rate. For this contract, it appears that the effort will be substantial, running 50 hours for the first session. Your firm bills at the rate of $100/hour and the overhead is expected to run a fixed $600 per session. The customer will pay you a flat fixed rate per session (Per Session Price.) If your profit markup is 20%, what will be the Total Price, the Per Session Price, and at what session will you break even?

After the above, answer the following three questions:

a. What is the Total Price? This is what you would charge the customer so that you can have your profit markup of 20% over all of your costs. To calculate this, first figure out your cost per each session, add them up, and then add your profit.

b. What is the Per Session Price? This is the revenue that the customer pays you each time you complete a session. It is calculated by dividing the Total Price by the number of sessions.

c. What is the Break Even Point? At the beginning, your cost per session is more than your revenue per session. Gradually, your cumulative revenue matches the cumulative cost, and eventually exceeds it so that you can end up with the desired profit. The break-even point is the session at which, for the first time, your revenue exceeds your cost.

Please show your excel calculations and formulas.

Please email Microsoft Excel Calculations to bundem2004@yahoo.com
Answer #1

Cumulative time taken for 8 sessions at a learning rate of 85% = 5.936 x 50 = 296.8 hours

total cost for 8 sessions = 296.8x100+600x8 = 34480

cost per session = 34480 /8 =4310

Total price of 8 sessions = 34480x1.2 =41376

Price per session = 41376 /8 =5172

Customer;s payment for each session = 5172

Cost per session for session 1 = 50x100+600 =5600

Cost per session for session 2 = 50x0.85x100 +600 = 4850

which is less than the monthly revenue. hene the breakeven will be achieved in session 2.